How much heat should be transferred when 20.1 g of hydrogen bromide is formed during the reaction of hydrogen gas with liquid bromine?
1
Expert's answer
2020-05-10T09:19:54-0400

Solution.

"m(HBr)=20.1g;"

"M(HBr)=81g\/mol;"

"\nu(HBr)=\dfrac{m(HBr)}{M(HBr)};"

"\nu(HBr)=\dfrac{20.1g}{81g/mol}=0.248mol;"

The enthalpy of formation of hydrogen bromide is equal to -34.1kJ/mol, that is, in the formation of one mole of hydrogen bromide, 34 kJ of heat is absorbed.

"1mol----34kJ;"

"0.248mol--xkJ;"

"x=\dfrac{0.248mol\sdot34kJ}{1mol}=8.437kJ;"


Answer: 8.437kJ of heat is absorbed.